# Tensor Decomposition Networks for Fast Machine Learning Interatomic Potential Computations

### v3

@@ Abstract
$\rm{SO}(3)$-equivariant networks are the dominant models for machine learning interatomic potentials (MLIPs). The key operation of such networks is the Clebsch-Gordan (CG) tensor product, which is computationally expensive. To accelerate the computation, we develop tensor decomposition networks (TDNs) as a class of approximately equivariant networks in which CG tensor products are replaced by low-rank tensor decompositions, such as the CANDECOMP/PARAFAC (CP) decomposition. With the CP decomposition, we prove (i) a uniform bound on the induced error of $\rm{SO}(3)$-equivariance, and (ii) the universality of approximating any equivariant bilinear map. To further reduce the number of parameters, we propose path-weight sharing that ties all multiplicity-space weights across the $\mathcal{O}(L^{3})$ CG paths into a single path without compromising equivariance, where $L$ is the maximum angular degree. The resulting layer acts as a plug-and-play replacement for tensor products in existing networks, and the computational complexity of tensor products is reduced from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We evaluate TDNs on PubChemQCR, a newly curated molecular relaxation dataset containing 105 million DFT-calculated snapshots. We also use existing datasets, including OC20, and OC22. Results show that TDNs achieve competitive performance with dramatic speedup in computations. Our code is publicly available as part of the AIRS library (https://github.com/divelab/AIRS/).
@@ 1 Introduction
Symmetry is a fundamental aspect of molecular and material systems (zhang2023artificial, ), making it a crucial consideration in developing machine learning interatomic potentials (MLIPs). Equivariant graph neural networks have emerged as dominant frameworks in this domain. Usually, equivariant models incorporate directional features and spherical harmonics to maintain equivariance under rotation symmetry (gasteiger2019directional, ; satorras2021n, ; schutt2021equivariant, ). Among these, tensor product (TP) operations play a central role in fusing equivariant features, providing a powerful mechanism for building expressive models that adhere to $\rm{SO}(3)$-equivariance (anderson2019cormorant, ; thomas2018tensor, ; liao2022equiformer, ; batatia2022mace, ).
However, the computational cost of tensor product operations grows rapidly with the maximum angular degree $L$, reaching $\mathcal{O}(L^{6})$ in conventional implementations. Recent efforts to mitigate this cost have focused on accelerating the TP operation or applying frame averaging (FA) to enforce equivariance (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ). While frame averaging is architecture-agnostic, it suffers from discontinuity issues (dym2024equivariant, ). On the other hand, TP acceleration techniques, such as SO(2)-based convolutions (passaro2023reducing, ; liao2024equiformerv, ) and fast spherical Fourier transformations (luoenabling, ), reduce complexity but are no longer the standard CG tensor product with the same expressivity. Thus, there remains a gap in developing a method that simultaneously reduces computational complexity and parameter count while maintaining similar accuracy and equivariance of CG tensor product.
In this work, we propose Tensor Decomposition Networks (TDNs), a new class of approximately equivariant networks that replace the standard CG tensor product with low-rank tensor decompositions based on CANDECOMP/PARAFAC (CP) decomposition. TDNs introduce a CP decomposition that provides error bounds on equivariance and universality, ensuring consistency under $\rm{SO}(3)$ transformations. Additionally, a path-weight sharing mechanism consolidates multiplicity-space weights across CG paths, significantly reducing the parameter count from $\mathcal{O}(cL^{3})$ to $\mathcal{O}(c)$ with $c$ the parameter count of weight of each path. The resulting layer has a close expressive power of conventional TP while lowering computational complexity from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We validate TDNs on a newly curated relaxation dataset with 105 million DFT-calculated molecular snapshots, along with the established OC20 and OC22 datasets, demonstrating competitive accuracy with substantial computational speedup.
@@ 2 Preliminaries and Related Work
Clebsch–Gordan (CG) tensor product is a fundamental operation widely used as the backbone for $\rm{SO}(3)$-equivariant neural network, enabling the fusion of feature fields at different angular degrees. We discuss the formal definition of CG tensor product with the maximum angular degree $L$. Consider two feature fields ${\bm{x}}=\bigoplus_{\ell_{1}=0}^{L}{\bm{x}}^{(\ell_{1})}$ and ${\bm{y}}=\bigoplus_{\ell_{2}=0}^{L}{\bm{y}}^{(\ell_{2})}$, where ${\bm{x}}^{(\ell_{1})},{\bm{y}}^{(\ell_{2})}$ are type-$\ell_{1}$ and type-$\ell_{2}$ irreducible representations (irreps) of $\rm{SO}(3)$. The CG tensor product fuses a pair of irreps ${\bm{x}}^{(\ell_{1})}$ and ${\bm{y}}^{(\ell_{2})}$ into every admissible output type $|\ell_{1}-\ell_{2}|\leq\ell_{3}\leq\ell_{1}+\ell_{2}$ via the CG coefficients $C^{\ell_{3},m_{3}}_{\ell_{1},m_{1},\ell_{2},m_{2}}$, defined as:
$({\bm{x}}^{(\ell_{1})}\!\otimes_{\!CG}\!{\bm{y}}^{(\ell_{2})})^{(\ell_{3})}_{m_{3}}=\sum_{m_{1}=-\ell_{1}}^{\ell_{1}}\sum_{m_{2}=-\ell_{2}}^{\ell_{2}}C^{\ell_{3},m_{3}}_{\ell_{1},m_{1},\ell_{2},m_{2}}\;{\bm{x}}^{(\ell_{1})}_{m_{1}}\,{\bm{y}}^{(\ell_{2})}_{m_{2}},\qquad-{\ell_{3}}\leq m_{3}\leq\ell_{3}.$
@@
${\bm{x}}\otimes_{\!CG}{\bm{y}}=\bigoplus_{\ell_{1}=0}^{L}\;\bigoplus_{\ell_{2}=0}^{L}\;\bigoplus_{\ell_{3}=|\ell_{1}-\ell_{2}|}^{\ell_{1}+\ell_{2}\leq L}({\bm{x}}^{(\ell_{1})}\!\otimes_{\!CG}\!{\bm{y}}^{(\ell_{2})})^{(\ell_{3})}.$
@@
However, the computational complexity of the CG tensor product scales as $\mathcal{O}(L^{6})$ as it involves $\mathcal{O}(L^{3})$ distinct paths and $\mathcal{O}(L^{3})$ operations per path. This significant cost poses a major bottleneck in practical implementations, especially when dealing with higher angular degrees. Computing the CG tensor product is very expensive due to the $\mathcal{O}(L^{6})$ computational complexity. To mitigate this computational challenge, inspired by tensor decomposition (kolda2009tensor, ) to low-rank tensors, we propose an efficient approximation based on tensor decomposition techniques, specifically the CP decomposition, to reduce the complexity while preserving approximate equivariance. The detailed formulation and implementation of the CP decomposition are presented in the following section.
@@
Invariant and Equivariant Models. Symmetry has been a widely discussed constraint when developing machine learning methods for predicting chemical properties of molecules. Invariant and equivariant graph models have been widely applied in these cases. Invariant models (schutt2018schnet, ; gasteiger2019directional, ; gasteiger2021gemnet, ; liu2022spherical, ; wang2022comenet, ) aim to consider the rotation invariant features such as pairwise distance as the input, and take use of these to predict final properties. Equivariant models (satorras2021n, ; schutt2021equivariant, ; deng2021vector, ; jing2021learning, ; tholke2022equivariant, ) further incorporate equivariant features such as pairwise directions and spherical harmonics into the model. These models are built with equivariant blocks to ensure that output features rotate consistently with any rotation applied to the input features, thereby maintaining equivariant symmetry.
@@
Tensor Product Acceleration. Among these equivariant networks, tensor product (anderson2019cormorant, ; thomas2018tensor, ; fuchs2020se, ; liao2022equiformer, ; batzner20223, ; batatia2022mace, ; unke2021se, ; yu2023efficient, ) is one of the most important components that fuse two equivariant features into one. It provides a powerful and expressive way (dym2020universality, ) to build equivariant networks, while the computational cost of TP is usually considerable. Therefore, there are several directions to accelerate the equivariant networks. First direction is to accelerate TP. eSCN (liao2024equiformerv, ; passaro2023reducing, ) proposes to reduce the SO(3) convolution into SO(2) for TP when one of the inputs of TP is spherical harmonics. Gaunt tensor product (luoenabling, ) makes use of fast spherical Fourier transformation to perform the TP. The other direction is to apply frame averaging (FA) (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ), which uses group elements from an equivariant set-value function called frame to transform the input data and subsequently the model’s output, enabling any models to obtain the desired symmetries. Although it is flexible and has no requirement for model architectures, it faces an unsolvable discontinuity problem (dym2024equivariant, ).
@@ 3.1 Tensor Product and Its Approximation
To reduce time complexity and parameter count of tensor product, we approximate the tensor $\bm{M}$} via CP decomposition (kolda2009tensor, ) of rank $R$. The CP decomposition writes the tensor product as sum of $R$ rank-1 tensors by decomposing the three-way tensor $\bm{M}$} such that
@@
where ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$, ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, and ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$ are factor matrices capturing the modes of the tensor. Substituting Eq.˜2 into Eq.˜1 gives
@@
For clarity, the above approximation for the order-two tensor product $m({\bm{x}},{\bm{y}})$ can be expressed in matrix form as:
@@
Next, we introduce the idea to make use of CP decomposition to accelerate the CG tensor product calculations. Specifically, the CG coefficient tensor is a three-way tensor concatenating CG coefficients ${\bm{C}}_{\ell_{1},\ell_{2}}^{\ell_{3}}$ of all admissible path $(\ell_{1},\ell_{2},\ell_{3})$ such that
${\bm{M}}=\bigoplus_{\ell_{1}=0}^{L}\bigoplus_{\ell_{2}=0}^{L}\bigoplus_{\ell_{3}=|\ell_{1}-\ell_{2}|}^{\ell_{1}+\ell_{2}\leq L}{\bm{C}}_{\ell_{1},\ell_{2}}^{\ell_{3}}.$
The key idea of CP decomposition is to break down the CG coefficient tensor $\bm{M}$} into low-rank matrices. We demonstrate a simple example of CP decomposition for TP with $\ell_{1},\ell_{2}=1$ and $\ell_{3}\in\{0,1,2\}$, as shown in Fig.˜1. Eq.˜3 allows for the simple batch-wise use of the Hadamard product with a global hyperparameter rank $R$ instead of computing tensor product per path. The higher $R$, the more accurate the tensor product result is. For acceleration, we do not need full rank $R$ and we can use a small $R$ while keeping a low approximation error. We show the rank-error curve in Fig.˜2.
@@ 3.2 Path-Weight Sharing Tensor Product
We first unify the multiplicities across all degrees $\ell$ by setting a common multiplicity $d$. Letting ${\bm{x}}^{(\ell)}\in\mathbb{R}^{d\times(2\ell+1)}$ denote the multiplicity-$d$ irrep of degree $\ell$, we concatenate the multiplicity axes into a single channel:
$\tilde{\bm{x}}=\operatorname{concat}_{\ell=0}^{L}{\bm{x}}^{(\ell)}\in\mathbb{R}^{d\times(L+1)^{2}}.$
@@
We further reduce the parameter count by applying a path-weight sharing scheme. Let ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell}$ denote the multiplicity-space weight matrix associated with the path $(\ell_{1},\ell_{2},\ell_{3})$, where $0\leq\ell_{1},\ell_{2},\ell_{3}\leq L$ and $c$ is the parameter count of weight of each path. Instead of assigning a unique matrix for each path, we set all such matrices equal to a single learnable parameter $\bm{W}$}, i.e. ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell}\equiv{\bm{W}}$ for every admissible path $(\ell_{1},\ell_{2},\ell_{3})$. This collapses the $\mathcal{O}(L^{3})$ distinct weight tensors of the naïve implementation into one, reducing the parameter count from $\mathcal{O}\bigl(cL^{3}\bigr)$ to $\mathcal{O}\bigl(c\bigr)$. Because the sharing operates exclusively on multiplicity indices, the irrep content of each block and hence full $\mathrm{SO}(3)$ equivariance is retained. The resulting layer therefore generalizes the classic CG tensor product while offering an order-of-magnitude reduction in parameters. We also extend this scheme to equivariant linear layers and use it in our main experiments.

### v4

@@ Abstract
$\rm{SO}(3)$-equivariant networks are the dominant models for machine learning interatomic potentials (MLIPs). The key operation of such networks is the Clebsch-Gordan (CG) tensor product, which is computationally expensive. To accelerate the computation, we develop tensor decomposition networks (TDNs) as a class of approximately equivariant networks in which CG tensor products are replaced by low-rank tensor decompositions, such as the CANDECOMP/PARAFAC (CP) decomposition. With the CP decomposition, we prove (i) a uniform bound on the induced error of $\rm{SO}(3)$-equivariance, and (ii) the universality of approximating any equivariant bilinear map. To further reduce the number of parameters, we propose path-weight sharing that ties all multiplicity-space weights across the $\mathcal{O}(L^{3})$ CG paths into a single shared parameter set without compromising equivariance, where $L$ is the maximum angular degree. The resulting layer acts as a plug-and-play replacement for tensor products in existing networks, and the computational complexity of tensor products is reduced from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We evaluate TDNs on PubChemQCR, a newly curated molecular relaxation dataset containing 105 million DFT-calculated snapshots. We also use existing datasets, including OC20, and OC22. Results show that TDNs achieve competitive performance with dramatic speedup in computations. Our code is publicly available as part of the AIRS library (https://github.com/divelab/AIRS/).
@@ 1 Introduction
Symmetry is a fundamental aspect of molecular and material systems (zhang2023artificial, ), making it a crucial consideration in developing machine learning interatomic potentials (MLIPs). Equivariant graph neural networks have emerged as dominant frameworks in this domain. Usually, equivariant models incorporate directional features and spherical harmonics to maintain equivariance under rotation symmetry (gasteiger2019directional, ; satorras2021n, ; schutt2021equivariant, ). Among these, tensor product (TP) operations play a central role in fusing equivariant features, providing a powerful mechanism for building expressive models that adhere to $\rm{SO}(3)$-equivariance (anderson2019cormorant, ; thomas2018tensor, ; liao2022equiformer, ; batatia2022mace, ; batzner20223, ).
However, the computational cost of tensor product operations grows rapidly with the maximum angular degree $L$, reaching $\mathcal{O}(L^{6})$ in conventional implementations. Recent efforts to mitigate this cost have focused on accelerating the TP operation or applying frame averaging (FA) to enforce equivariance (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ; tahmasebigeneralization, ). While frame averaging is architecture-agnostic, it suffers from discontinuity issues (dym2024equivariant, ). On the other hand, TP acceleration techniques, such as SO(2)-based convolutions (passaro2023reducing, ; liao2024equiformerv, ) and fast spherical Fourier transformations (luoenabling, ), reduce complexity but are no longer the standard CG tensor product with the same expressivity. Thus, there remains a gap in developing a method that simultaneously reduces computational complexity and parameter count while maintaining similar accuracy and equivariance of CG tensor product.
In this work, we propose Tensor Decomposition Networks (TDNs), a new class of approximately equivariant networks that replace the standard CG tensor product with low-rank tensor decompositions based on CANDECOMP/PARAFAC (CP) decomposition. TDNs introduce a CP decomposition that provides error bounds on equivariance and universality, ensuring consistency under $\rm{SO}(3)$ transformations. Additionally, a path-weight sharing mechanism consolidates multiplicity-space weights across CG paths, significantly reducing the parameter count from $\mathcal{O}(cL^{3})$ to $\mathcal{O}(c)$ with $c$ the parameter count of weight of each path. The resulting layer has expressive power close to conventional TP while lowering computational complexity from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We validate TDNs on a newly curated relaxation dataset with 105 million DFT-calculated molecular snapshots, along with the established OC20 and OC22 datasets, demonstrating competitive accuracy with substantial computational speedup.
@@ 2 Preliminaries and Related Work
Clebsch–Gordan (CG) tensor product is a fundamental operation widely used as the backbone for $\rm{SO}(3)$-equivariant neural network, enabling the fusion of feature fields at different angular degrees. We discuss the formal definition of CG tensor product with the maximum angular degree $L$. Consider two feature fields ${\bm{x}}=\bigoplus_{\ell_{1}=0}^{L}{\bm{x}}^{(\ell_{1})}$ and ${\bm{y}}=\bigoplus_{\ell_{2}=0}^{L}{\bm{y}}^{(\ell_{2})}$, where ${\bm{x}}^{(\ell_{1})},{\bm{y}}^{(\ell_{2})}$ are type-$\ell_{1}$ and type-$\ell_{2}$ irreducible representations (irreps) of $\rm{SO}(3)$. The CG tensor product fuses a pair of $\rm{SO}(3)$-representations ${\bm{x}}^{(\ell_{1})}$ and ${\bm{y}}^{(\ell_{2})}$ into every admissible output type $|\ell_{1}-\ell_{2}|\leq\ell_{3}\leq\ell_{1}+\ell_{2}$ via the CG coefficients $C^{\ell_{3},m_{3}}_{\ell_{1},m_{1},\ell_{2},m_{2}}$, defined as:
$({\bm{x}}^{(\ell_{1})}\!\otimes_{\!CG}\!{\bm{y}}^{(\ell_{2})})^{(\ell_{3})}_{m_{3}}=\sum_{m_{1}=-\ell_{1}}^{\ell_{1}}\sum_{m_{2}=-\ell_{2}}^{\ell_{2}}C^{\ell_{3},m_{3}}_{\ell_{1},m_{1},\ell_{2},m_{2}}\;{\bm{x}}^{(\ell_{1})}_{m_{1}}\,{\bm{y}}^{(\ell_{2})}_{m_{2}},\qquad-{\ell_{3}}\leq m_{3}\leq\ell_{3}.$ (1)
@@
${\bm{x}}\otimes_{\!CG}{\bm{y}}=\bigoplus_{\ell_{1}=0}^{L}\;\bigoplus_{\ell_{2}=0}^{L}\;\bigoplus_{\ell_{3}=|\ell_{1}-\ell_{2}|}^{\min(\ell_{1}+\ell_{2},\,L)}({\bm{x}}^{(\ell_{1})}\!\otimes_{\!CG}\!{\bm{y}}^{(\ell_{2})})^{(\ell_{3})}.$
@@
However, the computational complexity of the CG tensor product scales as $\mathcal{O}(L^{6})$ as it involves $\mathcal{O}(L^{3})$ distinct paths and $\mathcal{O}(L^{3})$ operations per path. This significant cost poses a major bottleneck in practical implementations, especially when dealing with higher angular degrees. To mitigate this computational challenge, inspired by tensor decomposition (kolda2009tensor, ) to low-rank tensors, we propose an efficient approximation based on tensor decomposition techniques, specifically the CP decomposition, to reduce the complexity while preserving approximate equivariance. The detailed formulation and implementation of the CP decomposition are presented in the following section.
@@
Invariant and Equivariant Models. Symmetry has been a widely discussed constraint when developing machine learning methods for predicting chemical properties of molecules. Invariant and equivariant graph models have been widely used in these cases. Invariant models (schutt2018schnet, ; gasteiger2019directional, ; gasteiger2021gemnet, ; liu2022spherical, ; wang2022comenet, ) aim to consider the rotation invariant features such as pairwise distance as the input, and make use of these to predict final properties. Equivariant models (satorras2021n, ; schutt2021equivariant, ; deng2021vector, ; jing2021learning, ; tholke2022equivariant, ; qu2024the, ) further incorporate equivariant features such as pairwise directions and spherical harmonics into the model. These models are built with equivariant blocks to ensure that output features rotate consistently with any rotation applied to the input features, thereby maintaining equivariant symmetry.
@@
Tensor Product Acceleration. Among these equivariant networks, tensor product (anderson2019cormorant, ; thomas2018tensor, ; fuchs2020se, ; liao2022equiformer, ; batzner20223, ; batatia2022mace, ; unke2021se, ; yu2023efficient, ) is one of the most important components that fuse two equivariant features into one. It provides a powerful and expressive way (dym2020universality, ) to build equivariant networks, while the computational cost of TP is usually considerable. Therefore, there are several directions to accelerate the equivariant networks. First direction is to accelerate TP. eSCN (passaro2023reducing, ; liao2024equiformerv, ; fu2025learning, ) proposes to reduce the SO(3) convolution into SO(2) for TP when one of the inputs of TP is spherical harmonics. Gaunt tensor product (luoenabling, ) makes use of fast spherical Fourier transformation to perform the TP. The other direction is to apply frame averaging (FA) (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ), which uses group elements from an equivariant set-valued function called frame to transform the input data and subsequently the model’s output, enabling any models to obtain the desired symmetries. Although it is flexible and has no requirement for model architectures, it faces an unsolvable discontinuity problem (dym2024equivariant, ).
@@ 3.1 Tensor Product and Its Approximation
To reduce time complexity and storage cost of tensor product, we approximate the tensor $\bm{M}$} via CP decomposition (kolda2009tensor, ) of rank $R$. The CP decomposition writes the tensor product as sum of $R$ rank-1 tensors by decomposing the three-way tensor $\bm{M}$} such that
@@
where ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$, ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, and ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$ are factor matrices capturing the modes of the tensor. Substituting Eq.˜3 into Eq.˜2 gives
@@
For clarity, the above approximation for the order-two tensor product $m({\bm{x}},{\bm{y}})$ can be expressed in matrix form as
@@
Next, we introduce the idea to make use of CP decomposition to accelerate the CG tensor product calculations. Specifically, the CG coefficient tensor is a three-way tensor concatenating CG coefficients $C_{\ell_{1},\ell_{2}}^{\ell_{3}}$ of all admissible path $(\ell_{1},\ell_{2},\ell_{3})$ such that
${\bm{M}}=\bigoplus_{\ell_{1}=0}^{L}\bigoplus_{\ell_{2}=0}^{L}\bigoplus_{\ell_{3}=|\ell_{1}-\ell_{2}|}^{\min(\ell_{1}+\ell_{2},\,L)}C_{\ell_{1},\ell_{2}}^{\ell_{3}}.$
The key idea of CP decomposition is to break down the CG coefficient tensor $\bm{M}$} into low-rank matrices. We demonstrate a simple example of CP decomposition for TP with $\ell_{1},\ell_{2}=1$ and $\ell_{3}\in\{0,1,2\}$, as shown in Fig.˜1. Eq.˜4 allows for the simple batch-wise use of the Hadamard product with a global hyperparameter rank $R$ instead of computing tensor product per path. The higher $R$, the more accurate the tensor product result is. For acceleration, we do not need full rank $R$ and we can use a small $R$ while keeping a low approximation error. We show the rank-error curve in Fig.˜2.
@@ 3.2 Path-Weight Sharing Tensor Product
We first unify the multiplicities across all degrees $\ell$ by setting a common multiplicity $k$. Letting ${\bm{x}}^{(\ell)}\in\mathbb{R}^{k\times(2\ell+1)}$ denote the multiplicity-$k$ irrep of degree $\ell$, we concatenate the irrep axes into a single channel:
$\tilde{\bm{x}}=\operatorname{concat}_{\ell=0}^{L}{\bm{x}}^{(\ell)}\in\mathbb{R}^{k\times(L+1)^{2}}.$
@@
We further reduce the parameter count by applying a path-weight sharing scheme. Let ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell_{3}}$ denote the multiplicity-space weight matrix associated with the path $(\ell_{1},\ell_{2},\ell_{3})$, where $0\leq\ell_{1},\ell_{2},\ell_{3}\leq L$ and $c$ is the parameter count of weight of each path. Instead of assigning a unique matrix for each path, we set all such matrices equal to a single learnable parameter $\bm{W}$}, i.e. ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell_{3}}\equiv{\bm{W}}$ for every admissible path $(\ell_{1},\ell_{2},\ell_{3})$. This collapses the $\mathcal{O}(L^{3})$ distinct weight tensors of the naïve implementation into one, reducing the parameter count from $\mathcal{O}\bigl(cL^{3}\bigr)$ to $\mathcal{O}\bigl(c\bigr)$. Because the sharing operates exclusively on multiplicity indices, the irrep content of each block and hence full $\mathrm{SO}(3)$ equivariance is retained. The resulting layer therefore retains the classic CG tensor product structure while offering an order-of-magnitude reduction in parameters. We also extend this scheme to equivariant linear layers and use it in our main experiments.